# USTC-TD: A Test Dataset and Benchmark for Image and Video Coding in 2020s

### v2

@@ IV-C Analysis of USTC-TD Dataset
In the above subsections, the construction of each dataset of USTC-TD is introduced, and we also point out some specific characteristics of the proposed dataset. To comprehensively verify the wide coverage of our proposed dataset for various content factors and qualitatively analyze the superiority of USTC-TD, we evaluate the USTC-TD on different image/video features and compare it with the previous image/video common test datasets (image datasets: Kodak[28], CLIC[30], Tecnick[29], video datasets: HEVC CTC[33], VVC CTC[34], MCL-JCV[32], UVG[31]). For analysis of image/video features, we select the spatial information (SI)[52], colorfulness (CF)[52], lightness information (LI)[53], and temporal information (TI)[31] to characterize each dataset along the dimensions of space, color, lightness, and temporal correlation, which are commonly used to evaluate the quality of dataset [54, 55, 31]. The definitions of the evaluation of these features can be found in [52, 31, 56, 53, 57], and the detailed calculation schemes are as follows:
@@
TABLE IX: Quantitative Results of the USTC-TD 2022 and 2023 Image Datasets. Note the Highest values are Represented in Red,
and the Lowest Values are Represented in Blue
@@
Note the Highest values are Represented in Red, and
the Lowest Values are Represented in Blue
@@
Video Sequence SI TI
USTC-Badminton 67.37 27.91
USTC-BasketballDrill 130.31 42.25
USTC-BasketballPass 94.63 46.92
USTC-BicycleDriving 38.66 52.70
USTC-Dancing 74.34 19.61
USTC-FourPeople 45.49 9.54
USTC-ParkWalking 72.48 40.22
USTC-Running 86.52 34.84
USTC-ShakingHands 100.67 44.11
USTC-Snooker 41.62 29.63
@@
Specifically, for the evaluation of USTC-TD image dataset, the scores of SI, LI, CF cooperate to evaluate its spatial, colorfulness and lightness diversity. Compared to the other image datasets[28, 30, 29], it exhibits SI scores ranging from 9 to 16, and is distinguished from other image datasets, as shown in Fig. 5 . The proposed image dataset incorporates more spatial diversity within the wide range of colorfulness diversity. In Fig. 5 , the proposed image dataset also shows a wide coverage of LI scores, ranging from 0.10 to 0.40, which aligns with the range of other image datasets and demonstrates that the proposed image dataset also exhibits excellent generalization of lightness diversity. For the evaluation of the USTC-TD video dataset, the scores of TI and SI cooperate to evaluate its spatial and temporal diversity, as shown in Fig. 5 . Compared to the other video datasets[33, 34, 32, 31], the USTC-TD exhibits a wide range of temporal variation with generalized spatial variation, ranging from 5 to 55. It further compensates for the absence of the 40 to 55 (higher) range in the temporal variation of other video datasets, which enables a wide coverage of temporal diversity for the comprehensive assessment of video compression-related algorithms.
In addition, we present the detailed variety distribution of these different features of USTC-TD image/video datasets, as shown in Table IX  and Table X . The higher/lower scores of different evaluative features are represented in different colors. Benefiting from the specific design of various content factors (environmental/imaging-related factors) of each image/video, we can find that the typical features of different content factors are evenly distributed in the captured image/video to ensure the diversity. For example, first, as mentioned in Section IV.A and Table VI , VII , IX , the setting of illumination (overexposure/underexposure) of USTC-2022-17, and USTC-2023-05, USTC-2023-09 enables the highest/lowest score of LI evaluative features, which demonstrates its specific design of imaging-related factors. Second, as mentioned in Section IV.B, Table VIII  and Table X  , the setting of complex motion (high-speed/tiny-speed moving objects) and lens (moving/fix) of USTC-BicycleDriving and USTC-FourPeople enables the higher/lower score of TI evaluative features, which demonstrates its specific design of environmental factors. We hope these specific design can help the researchers analyze and discover the bottlenecks of their research. The above analysis results and related codes are also open-sourced on the website.
Figure 6: Overall rate-distortion (RD) curves of advanced image compression schemes on PSNR and MS-SSIM metrics, the results are evaluated on USTC-TD image dataset 2022 and 2023.
@@ V Experiments
In this section, first, we present the experimental configurations employed for the evaluation of compression schemes. Second, we evaluate the classic standardized compression schemes and recent advanced image/video compression schemes on the proposed dataset under different metrics, and benchmark their performance on our proposed dataset. Third, we analyze the benchmarked performance, and further point out some limitations and inspirations among these advanced schemes to shed light on the future research.
@@ V-A Experimental Settings
In this subsection, first, we present the experimental settings of the evaluative compression schemes, including the selection of advanced image/video compression schemes and the training/testing configurations of these methods. Second, we introduce the evaluation metrics for these methods on our proposed dataset.
Figure 7: Overall rate-distortion (RD) curves of advanced video compression schemes on PSNR and MS-SSIM metrics, the results are evaluated on USTC-TD video dataset 2023.
V-A1 Selection of Evaluative Methods
For advanced image compression schemes, we select classic standardized schemes and advanced learned schemes. For traditional codecs, we select BPG [35] and H.266/VVC [2]. For learned image compression schemes, we select the Factorized Model[6], Hyperprior Model[6], Autoregressive Model[7], Cheng2020[8], iWave++[9], ELIC[10], and MLIC++[11]. For learned compression standardized schemes, we select the high-profile model of IEEE 1857.11[36] (iWave++).
For video compression schemes, we also select classic standardized schemes and advanced learned schemes. For standardized codecs, we select the H.265/HEVC[1], and H.266/VVC[2]. For learned video compression schemes, we select the DVC_Pro[12, 49], CANF-VC[16], DCVC[19], TCM-VC[20], DCVC-HEM[21], OOFE[22], SDD[23], VNVC[25], DCVC-DC[26], and DCVC-FM[27]. The detailed introduction and test instructions of these methods are mentioned in Section I of supplementary material.
@@ V-A2 Testing Configurations of Evaluative Traditional Image Compression Schemes
For testing, the officially released BPG software, VTM-17.0 (H.266/VVC reference software) are chosen. For BPG, the default configuration is used, and the internal color space is set to YUV444/420 for the testing of BPG and BPG444. For VTM-17.0, the encoder_intra_vtm configuration is used, and the internal color space is set to YUV444. For the different source formats of these testing datasets (USTC-TD and [28, 29, 30]), we convert them to the YUV444 color space for the input of VTM-17.0 by using the ffmpeg [39] tool (BT.601 conversion standard by default).
@@ V-A3 Testing Configurations of Evaluative Traditional Video Compression Schemes
For testing, the officially released HM-16.20 (HEVC reference software) and VTM-13.2 are chosen. For the setting of HM-16.20, the encoder_lowdelay_main_rext configuration is used. For the setting of VTM-13.2, the encoder_lowdelay_vtm configuration is used. The internal color space is set to YUV444. For the different source formats of these testing datasets (USTC-TD and [31, 32, 33, 34]), we also convert them to the YUV444 color space as the input of the above traditional codecs by using the ffmpeg [39] tool.
Figure 9: Specific rate-distortion (RD) curves of advanced video compression schemes on partial evaluative videos under PSNR metric.
@@ V-A4 Training and Testing Configurations of Evaluative Learned Image Compression Schemes
For training, these learned image compression models are optimized by mean squared error (MSE) or multi-scale structural similarity index measure (MS-SSIM), and the Flicker2W [60] dataset is used as the training dataset. These models are optimized by using the Adam Optimizer [61], with a batch size of 8 and a patch size of $256\times 256$. They are optimized for around 1.2 million iterations, starting with an initial learning rate of $10^{-4}$. The learning rate is reduced to $10^{-5}$ after 400 epochs and further down to $10^{-6}$ after 30 epochs. The setting of $\lambda$ is set to $\{0.001,0.004,0.024,0.080,0.200\}$ for iWave++, and $\{0.0018,0.0035,0.0067,0.0130,0.0250,0.0483\}$ for other schemes. For testing, the officially released model of MLIC++ and iWave++, and the reproduced model of ELIC are used. The model of other methods is provided by CompressAI [62].
@@ V-A5 Training and Testing Configurations of Evaluative Learned Video Compression Schemes
For training, these learned video compression models are mainly optimized by MSE or MS-SSIM, and the Viemo-90k[63] is used as the training set. For the testing of USTC-TD, the officially released models of these schemes are used. For the testing of the variable-rate model, the bitrate points are aligned to that of the traditional codec. For the testing of other models, we directly use the released model. For the different source formats of these testing datasets, we convert them from YUV444 to the RGB / YUV420 color space as the input of different learned video codecs by using the BT.709 conversion standard (adopted in JPEG AI[40]) / ffmpeg [39] tool.
V-A6 Evaluation Metrics
We use PSNR and MS-SSIM[37] to measure the quality of the coded frames in the comparison to the original frames. Bits per pixel (bpp) is used to measure the number of bits for encoding each pixel in each image or video frame. The Bjontegaard Delta bitrate (BD-rate)[64] is used to compare the performance of different compression schemes, where negative numbers indicate bitrate saving and positive numbers indicate bitrate increasing. For the evaluation of image and video compression schemes, the PSNR and MS-SSIM are both calculated and compared in RGB color space. For image, the conversion process of different color spaces is performed by using the ffmpeg tool [39]. For video, the conversion process is aligned to the setting mentioned in DCVC-DC[26] by using the BT.709 conversion standard.

### v3

@@ IV-C1 Characteristics of USTC-TD Image and Video datasets
To comprehensively verify the outstanding coverage of our proposed dataset for various content factors and qualitatively analyze the superiority of USTC-TD, we evaluate the USTC-TD on different image/video features and compare it with the previous image/video common test datasets (image datasets: Kodak[32], CLIC[34], Tecnick[33], video datasets: HEVC CTC[37], VVC CTC[38], MCL-JCV[36], UVG[35]). For analysis of image/video features, we select the spatial information (SI)[56], colorfulness (CF)[56], lightness information (LI)[57], and temporal information (TI)[35] to characterize each dataset along the dimensions of space, color, lightness, and temporal correlation, which are commonly used to evaluate the quality of dataset [58, 59, 35]. The definitions of the evaluation of these features can be found in [56, 35, 60, 57, 61], and the detailed ways are as follows:
@@
TABLE IX: Quantitative Results of the USTC-TD 2022 and 2023 Image Datasets. Note the Higher Scores are Represented in Red,
and the Lower Scores are Represented in Blue
@@
Note the Higher Scores are Represented in Red, and
the Lower Scores are Represented in Blue
@@
Video Sequence Short Videos Long Videos
SI TI SI TI
USTC-Badminton 67.37 27.91 71.35 27.94
USTC-BasketballDrill 130.31 42.25 131.66 44.19
USTC-BasketballPass 94.63 46.92 95.74 46.89
USTC-BicycleDriving 38.66 52.70 96.09 52.73
USTC-Dancing 74.34 19.61 7.22 26.16
USTC-FourPeople 45.49 9.54 45.77 9.59
USTC-ParkWalking 72.48 40.22 119.07 40.40
USTC-Running 86.52 34.84 89.02 34.97
USTC-ShakingHands 100.67 44.11 100.74 48.37
USTC-Snooker 41.62 29.63 50.01 35.85
@@
Specifically, for the evaluation of USTC-TD image dataset, the scores of SI, LI, CF cooperate to evaluate its spatial, colorfulness, and lightness diversity. Compared to the other image datasets[32, 34, 33], it exhibits SI scores ranging from 9 to 16, and is distinguished from other image datasets, as shown in Fig. 5 . The proposed image dataset incorporates more spatial diversity within the wide range of colorfulness diversity. In Fig. 5 , the proposed image dataset also shows a wide coverage of LI scores, ranging from 0.10 to 0.40, which aligns with the range of other image datasets and demonstrates that the proposed image dataset also exhibits excellent generalization of lightness diversity. For the evaluation of USTC-TD video dataset (short/long setting), the scores of TI and SI cooperate to evaluate its spatial and temporal diversity, as shown in Fig. 5 . Compared to the other video datasets[37, 38, 36, 35], the proposed video dataset incorporates more temporal diversity within the excellent generalization of spatial diversity. For the short setting, it exhibits a wide range of temporal variation, ranging from 5 to 55. It compensates for the absence of the 40 to 55 (higher) range in the temporal variation of other video datasets, which enables an excellent coverage of temporal diversity with the collaboration of other datasets. For the long setting, it further extends the coverage of temporal variation, enabling a more robust and comprehensive assessment of video compression schemes. The above analysis results and related codes are open-sourced.
IV-C2 Driving Factors Behind the Benefit of USTC-TD
Beyond the above feature analysis, we present the detailed variety distribution of these different features of USTC-TD image/video datasets, as shown in Table IX  and X , the detailed scores of different evaluative features are represented with the annotation of different colors. Here we deeply discuss the driving factors behind the benefit of USTC-TD image and video datasets.
TABLE XI: Three Recommended Presets of the Existing Image/Video Datasets [32, 33, 34, 35, 36, 37, 38] and the Proposed USTC-TD
for the Practical Evaluation of Image/Video Compression Schemes
Preset Image Dataset Video Dataset Characteristic Purpose
Challenging USTC-TD USTC-TD Challenging Content Developing Advanced Coding Algorithms
Desirable USTC-TD, Tecnick[33] USTC-TD, MCL-JCV[36], HEVC CTC[37] Wider Coverage Obtaining Reliable Estimation of Coding Performance
Ideal
USTC-TD, CLIC[34],
Tecnick[33], Kodak[32]
USTC-TD, UVG[35], MCL-JCV[36],
HEVC CTC[37], VVC CTC[38]
Sufficient Coverage
Evaluating Algorithm Robustness
and Generalization Ability
For the benefit of USTC-TD image dataset, it compensates for the absence of the higher range of spatial variation within the generalized colorfulness, lightness variation and adds the mixture content diversity, which distinctly differentiates it from existing image datasets[32, 33, 34]. Compared with the previous image datasets, the driving factors of the benefit mainly come from two aspects: the specific design of various content factors (environmental/imaging-related factors), and the mixture of content diversity (spatial, lightness, colorfulness diversity), which contribute to maximizing the coverage. In detail, first, the samples with higher spatial diversity exhibit the characteristics of complex content factors. For example, the samples with SI scores higher than 13 of Fig. 5  have the above aspect, such as the USTC-TD-2022-01, 14, 18, USTC-TD-2023-04, 05, 08. According to the Table VIII  and VIII , these samples own the specific design of complex environmental or imaging factors, such as the USTC-TD-2022-01, 18, USTC-TD-2023-04 with the typical texture (geometric), or USTC-TD-2022-14/USTC-TD-2023-05 with extreme illumination (underexposure/overexposure), or USTC-TD-2023-08 with the special captured view (overhead level). Second, the samples with mixture diversity enable the outstanding coverage of different features, such as the USTC-TD-2022-07, 11, USTC-TD-2023-05, 15. As shown in Table IX , these samples own the mixture of different kinds of content diversity, such as the USTC-TD-2022-07 with higher spatial and lightness diversity (higher SI, LI scores), the USTC-TD-2023-05 with various diversity (higher SI, Lower CF, LI scores). The mutual promotion of different components of content factors has further contributed to the maximization of coverage.
For the benefit of USTC-TD video dataset, it compensates for the absence of the higher range of temporal variation within the generalized spatial variation, which distinctly differentiates it from existing video datasets[37, 38, 36, 35] . Compared with the previous video datasets, the driving factors of the benefit directly come from the augmentation and extension of motion diversity. For example, according to Table VIII and X , the samples with TI scores higher than 40 of Fig. 5  exhibit the characteristics of diverse motion factors, such as USTC-BicycleDriving with high-speed moving objects and scene change, USTC-ShakingHands with precise limb motion, USTC-BaskerballDrill with abundant athletes’ motion and rich lens moving. Compared with the characteristics of existing datasets, the above different kinds of motion factors further compensate for the limitation of temporal correlation, and contribute to the robust evaluation and the maximization of temporal relation-related property coverage.
IV-C3 Discussion of Practical Utilization and Application of the Existing Compression Datasets and USTC-TD
Beyond the feature and benefit analysis of USTC-TD, it has been verified that the USTC-TD compensates for the shortcomings of existing image/video datasets, and its collaboration with existing datasets can achieve wider coverage of different image/video features. Although the collaboration of these datasets ensures the robust evaluation of advanced compression schemes, the evaluation of all image/video datasets will cause heavy time/computational complexity in practical use. Here we discuss their desirable collaboration for future compression research, and further put forward the different recommended presets of USTC-TD and existing image/video datasets with the consideration of practical utilization.
To achieve an efficient evaluation for future research, we discuss the desirable collaboration of existing image/video datasets and USTC-TD from two perspectives: challenge and coverage, and present three presets of these image and video datasets to support them with the consideration of different uses, including challenging, desirable, and ideal presets, as shown in Table XI . For the challenge, in the development of compression standardization, the coding efficiency on existing test content has gradually reached a bottleneck, such as the performance of hand-crafted intra and inter prediction, making it increasingly difficult to reflect shortcomings and guide future improvements. Thus, incorporating challenging content is essential, as it serves as a continuous driving force for the improvement of next-generation coding standards. For challenging preset, typical challenging samples are considered for developing advanced coding algorithms. Correspondingly, USTC-TD samples with the highest spatial/temporal diversity (scores) are chosen for evaluation. For coverage, building upon the challenging preset, a broader coverage of diverse content is desirable for the consistent evaluation of algorithm performance. By incorporating a wider range of different kinds of content features, test datasets can effectively evaluate the codec’s performance in various scenarios, ensuring its reliability in practical use. For desirable preset, the samples with the highest coverage of different kinds of content diversity (spatial, colorfulness, lightness for image samples, temporal, spatial for video samples) are considered for obtaining reliable estimation of coding performance, such as the collaboration of USTC-TD image dataset and Tecnick [33], USTC-TD video dataset, MCL-JCV[36], and HEVC CTC[37]. These collaborations cover the majority of samples with different kinds of content diversity, with an appropriate number of representative samples supporting comprehensive evaluation in practical use. For the ideal preset, given sufficient time/resources, all datasets are ideally tested to ensure a comprehensive evaluation of algorithm robustness and generalization, supporting its use in rigorous benchmarking and evaluation.
@@ V Experiments
In this section, first, we present the experimental configurations employed for the evaluation of compression schemes. Second, we evaluate the classic standardized compression schemes and recent advanced image/video compression schemes on the proposed dataset under different metrics, and benchmark their performance on our proposed dataset. Third, we analyze the benchmarked performance, and further point out some limitations and inspirations among these advanced image/video schemes to shed light on future research.
@@ V-A Experimental Settings
In this subsection, first, we present the experimental settings of the evaluative compression schemes, including the selection of advanced image/video compression schemes and the training/testing configurations of these compression schemes. Second, we introduce the evaluative quality metrics for these schemes on our proposed dataset.
V-A1 Selection of Evaluative Compression Schemes
For advanced image compression schemes, we select classic standardized schemes and advanced learned schemes. For traditional codecs, we select BPG [^6] Available online at  https://bellard.org/bpg., and H.266/VVC [2]. For learned image compression schemes, we select the Factorized Model[6], Hyperprior Model[6], Autoregressive Model[7], Cheng2020[8], iWave++[9], ELIC[10], and MLIC++[11]. For learned compression standardized schemes, we select the high-profile model of IEEE 1857.11777Available online at  https://sagroups.ieee.org/fvc/. (iWave++).
For video compression schemes, we also select classic standardized schemes and advanced learned schemes. For standardized codecs, we select the H.265/HEVC[1], H.266/VVC[2], AV1[4], and AV2 [5]. For learned video compression schemes, we select the DVC_Pro[12, 53], CANF-VC[16], DCVC[19], TCM-VC[20], DCVC-HEM[21], OOFE[22], SDD[23], VNVC[25], DCVC-DC[26], and DCVC-FM[27]. The detailed introduction and test instructions of these methods are mentioned in Section I of supplementary material.
@@ V-A2 Testing Configurations of Evaluative Traditional Image Compression Schemes
For testing, the officially released BPG software, VTM-17.0 (H.266/VVC reference software) are chosen. For BPG, the default configuration is used, and the internal color space is set to YUV420/444 for the testing of BPG and BPG444. For VTM-17.0, the encoder_intra_vtm configuration is used, and the internal color space is set to YUV444. For the different source formats of these testing datasets (USTC-TD and [32, 33, 34]), we convert them to the YUV444 color space for the input of VTM-17.0 by using the default ffmpeg tool (BT.601 conversion standard).
@@ V-A3 Testing Configurations of Evaluative Traditional Video Compression Schemes
For testing, the officially released HM-16.20 (H.265/HEVC reference software), VTM-13.2 (H.266/VVC reference software), AV1-3.11.0, and AV2-7.0.0 are chosen. For the setting of HM-16.20, the encoder_lowdelay_main_rext configuration is used. For the setting of VTM-13.2, the encoder_lowdelay_vtm configuration is used. For the different source formats of these testing datasets (USTC-TD and [35, 36, 37, 38]), we convert them to the YUV444 color space as the input of the above traditional codecs by using the default ffmpeg tool, the conversion step is aligned to the setting mentioned in DCVC-DC[26].
Figure 6: Overall rate-distortion (RD) curves of advanced image compression schemes on different metrics. From left to right, the results are evaluated by PSNR, MS-SSIM, VMAF (MSE model), and VMAF (MS-SSIM model) metrics on USTC-TD image dataset 2022 and 2023.
(a) Short Setting, Intra Period = 32
(b) Long Setting, Intra Period = 32
(c) Long Setting, Intra Period = -1
Figure 7: Overall rate-distortion (RD) curves of advanced video compression schemes on different metrics. From left to right, the results are evaluated by PSNR, MS-SSIM, VMAF (MSE model), and VMAF (MS-SSIM model) metrics on different settings of USTC-TD video dataset.
@@ V-A4 Training and Testing Configurations of Evaluative Learned Image Compression Schemes
For training, these learned image compression models are optimized by mean squared error (MSE) or multi-scale structural similarity index measure (MS-SSIM), and the Flicker2W [64] dataset is used as the training dataset. These models are optimized by using the Adam Optimizer [65], with a batch size of 8 and a patch size of $256\times 256$. They are optimized for around 1.2 million iterations, starting with an initial learning rate of $10^{-4}$. The learning rate is reduced to $10^{-5}$ after 400 epochs and further down to $10^{-6}$ after 30 epochs. The setting of $\lambda$ is set to $\{0.001,0.004,0.024,0.080,0.200\}$ for iWave++, and $\{0.0018,0.0035,0.0067,0.0130,0.0250,0.0483\}$ for other schemes. For testing, the officially released model of MLIC++/iWave++, the reproduced model of ELIC are used. The model of other schemes is provided by CompressAI [66].
@@ V-A5 Training and Testing Configurations of Evaluative Learned Video Compression Schemes
For training, these learned video compression models are mainly optimized by MSE or MS-SSIM, and the Viemo-90k[67] is used as the training dataset. For the testing of USTC-TD, the officially released models of these schemes are used. For the testing of the variable-rate model, the bitrate points are aligned to that of the traditional codec. For the testing of other models, we directly use the released model. For the different formats of these testing datasets, we convert them to the RGB color space as the input of different learned video codecs by using the default ffmpeg tool (BT.601), the conversion step is aligned to the setting mentioned in DCVC-DC[26].
V-A6 Testing Configurations of Subjective Quality Evaluation of Image and Video Compression Schemes
For the testing of traditional compression model, we directly use their officially released standardized model, these hand-crafted models are designed and optimized for PSNR metric. For the testing of learned compression model, we use the officially released MS-SSIM model of the advanced learned image and video compression schemes, these trained models are optimized for MS-SSIM metric. Some image compression schemes [9, 10] without the open-sourced MS-SSIM model are skipped. For the testing of video compression model, the long USTC-TD video dataset is used. To efficiently conduct the subjective test, we select the compressed images/video results at the intermediate bitrate point among all test bitrate points. For the evaluation of image compression schemes, since most learned approaches do not support the variable bitrate models, we align the target bitrate point of all approaches to approximately 0.12 bits per pixel. For the evaluation of video compression schemes, we align the target bitrate point of all approaches to approximately 0.06 bits per pixel.
V-A7 Evaluative Metrics
For the objective quality evaluation, PSNR, MS-SSIM[39], and VMAF[40] are used to measure the quality of the coded frames in comparison to the original frames. Bits per pixel (bpp) is used to measure the number of bits for encoding each pixel in each image/video frame. The Bjontegaard Delta bitrate (BD-rate)[68] is used to compare the performance of different compression schemes, where negative numbers indicate bitrate saving and positive numbers indicate bitrate increasing. For the evaluation of image/video compression schemes, PSNR and MS-SSIM are calculated and compared in RGB color space, VMAF is calculated and compared in YUV color space. The conversion process of different color spaces is performed by using the default ffmpeg tool (BT.601). For PSNR and MS-SSIM tests, the MSE model of these schemes is used for PSNR, while the MS-SSIM model of these schemes is used for MS-SSIM. For VMAF test, the MSE and MS-SSIM models are all used for VMAF test (VMAF MSE/MS-SSIM model for short). Notably, for some test schemes without the MS-SSIM model, their MSE model is used for the MS-SSIM test, and the VMAF (MS-SSIM model) test is skipped.
[…]
$MOS=\frac{1}{M}\sum_{i=1}^{M}\left(\frac{1}{N}\sum_{j=1}^{N}S_{i,j}\right)$ (5)
where $M$ is the total number of test samples, $N$ is the number of participants, and $S_{i,j}$ represents the rating given by the $j$-th participant for the $i$-th sample. In our setting, 50 participants are selected for the MOS test, with testers from USTC and online sources, ensuring diversity in user experience and perception.